# Selective Mixup for Debiasing Question Selection in Computerized Adaptive Testing

### v1

@@ 3.2. Notation and Problem Setting
We divide the examinees in $\mathcal{D}_{m}$ into three attributes: $\mathcal{D}_{m}^{\text{A}},\ \mathcal{D}_{m}^{\text{B}},\ \mathcal{D}_{m}^{\text{C}}$, based on their proportion of correct responses. These groups represent different meta set distributions. Let $\mathcal{A}=\{\text{A},\text{B},\text{C}\}$ denote the attribute categories.
  • A: the proportion of correct responses is in the range $[0,0.4]$, where incorrect responses dominate.
  • B: the proportion of correct responses falls within $(0.4,0.6]$, indicating a roughly balanced distribution.
  • C: the proportion of correct responses is in the range $(0.6,1.0]$, where correct responses dominate.
We also define the label set as $\mathcal{Y}=\{0,1\}$, where 0 denotes an incorrect response and $1$ denotes a correct response. A group is defined as a unique combination of an attribute $a\in\mathcal{A}$ and a label $y\in\mathcal{Y}$. Based on this group definition, we further categorize the interactions into three types:
  • Bias-aligned samples: Interactions that conform to the dominant attribute pattern and are typically easy to learn. This includes interactions from $\mathcal{D}_{m}^{\text{A}}$ with label 0, and from $\mathcal{D}_{m}^{\text{C}}$ with label $1$.
  • Unbiased samples: Interactions from $\mathcal{D}_{m}^{\text{B}}$, where responses are balanced and do not exhibit strong bias. These are considered neutral.
  • Bias-conflicting samples: Interactions that go against the dominant attribute pattern and are typically hard to learn. This includes interactions from $\mathcal{D}_{m}^{\text{A}}$ with label $1$, and from $\mathcal{D}_{m}^{\text{C}}$ with label 0.
In CAT, the examinee’s true proficiency $\theta_{i}^{*}$ is assumed constant, and traditional methods aim to make the final estimate $\theta_{i}^{T}$ closely approximate it. This work instead focuses on mitigating selection bias in the data-driven Question Selection Module to produce accurate, unbiased proficiency estimates and maintain robustness under distributional shifts.
@@ 4. Method
This section presents our approach to mitigating bias in data-driven CAT. The method addresses selection bias in the meta set through a selective mixup strategy guided by examinee attributes. By generating augmented samples from balanced response profiles, it enriches underrepresented cases and helps the Question Selection Module learn smoother, less biased decision boundaries. The approach is simple, effective, and can be readily integrated into data-driven question selection in CAT.
@@ 4.1. Overview
Our method mitigates selection bias in CAT through a two-stage framework: Cross-Attribute Examinee Retrieval and Selective Mixup Regularization, as illustrated in Figure 3. In the first stage, for each examinee in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$, the most similar counterpart is retrieved from $\mathcal{D}_{m}^{\text{B}}$ based on response patterns. In the second stage, the focus is on bias-conflicting samples, such as correct responses from $\mathcal{D}_{m}^{\text{A}}$ or incorrect responses from $\mathcal{D}_{m}^{\text{C}}$. To enrich these underrepresented cases, each examinee from $\mathcal{D}_{m}^{\text{B}}$ is paired with a counterpart from $\mathcal{D}_{m}^{\text{A}}$ or $\mathcal{D}_{m}^{\text{C}}$, and mixup is applied only when their interaction labels are identical. The final training objective combines the empirical loss on the original data with additional regularization from synthetic samples, thereby improving the robustness and generalization capability of the Question Selection Module under biased meta distributions.
Figure 3. An overview of our debiasing framework for Computerized Adaptive Testing. The method consists of two stages: (1) Cross-Attribute Examinee Retrieval, where we retrieve midperformers from $\mathcal{D}_{m}^{\text{B}}$ as neutral references for examinees in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$ based on proficiency similarity; and (2) Selective Mixup-based Regularization, where mixup is applied between retrieved midperformers and their matched counterparts (with the same label) to generate synthetic training samples. This augmentation improves the robustness of the Question Selection Module by reducing reliance on bias-aligned patterns.
@@ 4.2. Cross-Attribute Examinee Retrieval
In real-world CAT scenarios, the training data often exhibits significant selection bias: examinees in $\mathcal{D}_{m}^{\text{A}}$ are more likely to respond incorrectly, while those in $\mathcal{D}_{m}^{\text{C}}$ tend to answer most questions correctly. As a result, the training set is dominated by bias-aligned interactions, such as incorrect responses from $\mathcal{D}_{m}^{\text{A}}$ and correct responses from $\mathcal{D}_{m}^{\text{C}}$, which are easy for the Question Selection Module to fit but may lead to shortcut learning and reduced robustness. In contrast, bias-conflicting samples, including correct responses from $\mathcal{D}_{m}^{\text{A}}$ and incorrect responses from $\mathcal{D}_{m}^{\text{C}}$, are much rarer and more challenging to learn, yet they are essential for capturing deeper examinee–question relationships and ensuring fairness and generalization. To alleviate this imbalance, we leverage examinees from $\mathcal{D}_{m}^{\text{B}}$, whose balanced response patterns resemble an unbiased distribution and provide reliable references for augmenting the underrepresented bias-conflicting cases, which naturally motivates our selective mixup design.
Concretely, for each examinee $e_{a}$ with attribute $a\in\{\text{A},\text{C}\}$, we retrieve the most similar examinee $e_{b}$ with attribute $b\in\text{B}$ based on their cognitive proficiency vectors $\boldsymbol{\theta}$, estimated by the Cognitive Diagnosis Module. We compute the similarity via the $L_{2}$ norm:
(9) $\text{similarity}_{i,j}=\left\|\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{j}\right\|_{2},$
where $\boldsymbol{\theta}_{i}$ and $\boldsymbol{\theta}_{j}$ represent the latent knowledge proficiency vectors of examinees $i$ and $j$, respectively.
This retrieval process serves two purposes: first, it anchors bias-conflicting interactions with semantically similar yet statistically more balanced samples; second, it establishes the basis for subsequent augmentation. By aligning examinees in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$ with appropriate references from $\mathcal{D}_{m}^{\text{B}}$, the method effectively bridges biased extremes and enables the model to learn smoother decision boundaries in the next stage.
@@ 4.3. Selective Mixup-based Regularization
To further mitigate selection bias, we propose a selective mixup augmentation strategy (zhang2018mixup). Unlike conventional training that heavily depends on naturally occurring samples, mixup introduces synthetic interactions by interpolating instances, which substantially increases the diversity of bias-conflicting data. This is particularly important in CAT, where the scarcity of such cases otherwise causes the Question Selection Module to overfit to bias-aligned patterns and overlook minority interactions. Mixup has several key advantages in this context: it smooths decision boundaries, reduces overfitting, and improves generalization when the test distribution differs from the training distribution. Moreover, by diluting the dominance of spurious correlations embedded in the original data, mixup encourages the model to learn more balanced and robust representations. Through this targeted augmentation, the Question Selection Module is guided toward more reliable and unbiased predictions, ultimately enhancing its robustness and generalization under distribution shifts and ensuring more consistent performance across diverse examinees.
@@
Given $q_{i}\in{\mathcal{D}_{m}^{\text{A}},\mathcal{D}_{m}^{\text{C}}}$, its most similar counterpart in terms of proficiency is first identified from $\mathcal{D}_{m}^{\text{B}}$ in the previous stage. From this matched examinee in $\mathcal{D}_{m}^{\text{B}}$, an interaction $q_{j}$ with the same true label $y$ is then sampled and paired with $q_{i}$, upon which synthetic bias-conflicting samples are generated through convex combination. With such paired questions $(q_{i},q_{j})$, we further interpolate their associated parameters to construct fake questions. The representation mixup is as follows:
(10) $\boldsymbol{Q}_{i,j}^{\text{fake}}$ $=\lambda\boldsymbol{Q}_{i}+(1-\lambda)\boldsymbol{Q}_{j},$
(11) $\boldsymbol{h}_{i,j}^{\text{diff}}$ $=\lambda\boldsymbol{h}^{\text{diff}}_{i}+(1-\lambda)\boldsymbol{h}_{j}^{\text{diff}},$
(12) $h_{i,j}^{\text{disc}}$ $=\lambda h_{i}^{\text{disc}}+(1-\lambda)h_{j}^{\text{disc}}.$
@@
The synthesized features are then fed into the CDM, and the corresponding response probability is computed as $M(q_{i,j}^{\text{fake}};\boldsymbol{\theta},\zeta_{q_{i,j}^{\text{fake}}},\zeta_{n})$. Since the mixup is performed between samples sharing the same label, the synthetic sample inherits the same label as the originals. The loss for these augmented samples is therefore defined as:
(13) $\mathcal{L}_{\text{fake}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{q_{i,j}\in\mathcal{D}_{m}}\ell\!\left(y_{i,q_{i,j}^{\text{fake}}},M\!\left(q_{i,j}^{\text{fake}};\boldsymbol{\theta}_{i}^{*},\zeta_{q_{i,j}^{\text{fake}}},\zeta_{n}\right)\right).$
@@
The final outer-level optimization objective can be rewritten as the combination of empirical risk minimization and synthetic sample regularization:
(14) $\mathcal{L}_{\text{emp}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{q_{i,j}\in\mathcal{D}_{m}}\ell\!\left(y_{i,q_{i,j}},M\!\left(q_{i,j};\boldsymbol{\theta}_{i}^{*},\zeta_{q_{i,j}},\zeta_{n}\right)\right),$
(15) $\min_{\zeta_{s}}\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{emp}}+\omega\mathcal{L}_{\text{fake}}.$
where $\omega$ controls the trade-off between real and synthetic samples.
The dual-objective design encourages the model to maintain strong predictive performance on authentic data while simultaneously enhancing its robustness to bias-conflicting cases. The outer-level optimization updates the Question Selection Module parameters $\zeta_{s}$ via gradient descent on the combined loss function. A full overview of the training process, including mixup-based augmentation and bias-aware optimization steps, is provided in Algorithm 1, which outlines how the proposed strategy is integrated into the outer-loop optimization to enhance robustness and mitigate selection bias.
Algorithm 1 Debiasing in CAT Training Process
1: Initialize parameters $\boldsymbol{\theta}$ for the Cognitive Diagnosis Module using pre-trained parameters $(\zeta_{q},\zeta_{n})$, and initialize $\zeta_{s}$ for the Question Selection Module. Set learning rates $\alpha$, $\beta$, number of gradient descent steps $K$ for inner-level optimization, and number of test steps $T$.
@@
9:   Apply selective mixup augmentation using Eqs. (10) and (13).
@@
13:return Estimated proficiencies $\{\boldsymbol{\theta}_{i}^{*}\}_{i=1}^{N}$ and debiased Question Selection Module $\zeta_{s}$.
@@
By interpolating between samples from attribute A or C and their corresponding references from attribute B, our approach introduces balanced, label-consistent synthetic data into the training process. This prevents the Question Selection Module from overfitting to dominant response patterns and encourages it to capture more nuanced relationships between question difficulty and proficiency. In effect, the model becomes less reliant on biased historical distributions, improving fairness and robustness across diverse examinees.
@@ 5. Experiments
In this section, we first introduce the dataset and experimental setup. Then, we conduct extensive experiments on real-world datasets. The code in our experiments are available at: https://anonymous.4open.science/r/MDCAT.
@@ 5.1. Experimental Settings
In this section, we introduce the datasets, the selected baselines, and the application of CAT.
DataSet Description. We conducted experiments on two widely used educational datasets. The ASSIST0910 dataset [^2] https://sites.google.com/site/assistmentsdata/home/assistment2009-2010-data contains student practice logs from the ASSISTments online tutoring system, focusing on mathematics and knowledge concept problems. The NIPS-EDU dataset [^3] https://eedi.com/projects/neurips-education-challenge, released as part of the NeurIPS 2020 Education Challenge, includes student responses collected from the Eedi learning platform. We used the interaction data from Task 3 and Task 4. For the ASSIST0910 dataset, we first removed records containing missing values and subsequently excluded examinees with fewer than 40 interactions. The dataset statistics, including the number of examinees, questions, concepts, and interactions, are presented in Table 1.
@@
Partition. We split the students into training (60%), validation (20%), and testing (20%) sets. For evaluation, we conduct 5 rounds of validation on each dataset. In each round, we further partition the questions answered by each student into a support set ($\mathcal{D}_{u}$, 80%) and a meta set ($\mathcal{D}_{m}$, 20%). To prevent overfitting, these partitions are randomly regenerated at the beginning of each training epoch.
Evaluation.To assess model performance, we evaluate all methods based on their ability to predict binary-valued responses in the meta set $\mathcal{D}_{m}$, using two metrics: the worst-group accuracy (Worst) and the average accuracy across all samples (Avg). All models are implemented in PyTorch and trained on an NVIDIA 3090 GPU.
Baseline Methods. For Cognitive Diagnosis Modules, we adopt the 1PL IRT model (embretson2013irt) and NCDM (wang2020ncd), both of which are implemented in the EduStudio [^4] https://edustudio.ai/ framework (wu2025edustudio). For CAT frameworks, we focus on representative data-driven approaches that formulate CAT as a bi-level optimization problem and solve it via meta-learning. Specifically, we include BOBCAT (ghosh2021bobcat) and UATS (yu2024ucats), both of which jointly optimize the Question Selection Module and the student model. BOBCAT formulates the problem as a bi-level optimization task with a learnable Question Selection Module, while UATS employs a differentiable hierarchical optimization framework to enhance both efficiency and adaptability in question selection. To better analyze bias in Question Selection Module, we freeze the question-related parameters of the Cognitive Diagnosis Module during outer-loop optimization and instead initialize them with pre-trained values.
To validate the effectiveness of our method, we compare it with the following debiasing baselines, all implemented within the same bi-level training framework and adapted to the context of bias mitigation in CAT.
ERM (Empirical Risk Minimization) minimizes the average prediction loss over the entire meta set, without explicitly addressing distributional bias. It serves as a standard baseline but tends to overfit to dominant bias-aligned patterns, failing to generalize to bias-conflicting samples.
IRM (arjovsky2019irm) (Invariant Risk Minimization) encourages the model to learn feature representations that support consistent predictions across different environments. In our setting, we simulate multiple training environments by sampling examinees with different attribute values (i.e., A, B, C) and ensure that the Question Selection Module performs well across these diverse distributions.
GroupDRO (sagawa2019groupDRO) (Group Distributionally Robust Optimization) focuses on improving the worst-case performance across predefined groups. We optimize the Question Selection Module by minimizing the maximum group-wise loss, thus preventing performance collapse on underrepresented or difficult subpopulations.
Reweight (japkowicz2000reweight) directly addresses imbalance by assigning larger weights to bias-conflicting samples during training. This strategy guides the model to better capture underrepresented patterns and reduces reliance on shortcut signals in bias-aligned data.
Parameter Settings. In NCDM, 128 and 64 denote the number of neurons in the two hidden layers, which control the capacity and complexity of the network. For the $\alpha$ parameter in the Beta distribution, we set the scale to $\{0.2,0.4,0.6,0.8,1.0\}$. Similarly, for the mixing coefficient $\omega$ in the mixup loss, we use values from the same set: $\{0.2,0.4,0.6,0.8,1.0\}$. To approximate each student’s ability, we perform inner optimization by updating the student’s latent ability vector $K$ times, with $K=5$ in our experiments.
Table 2. Performance of Different CAT Models under IRT-based and NCDM-based Cognitive Diagnosis with Various Debiasing Methods on the Unbiased ASSIST0910 and NIPS-EDU Datasets. Here, Worst denotes worst‑group accuracy and Avg. denotes overall average accuracy.
@@
Metrics@10 Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg.
@@
Ours 0.4489 0.6320 0.4160 0.6149 0.4426 0.6117 0.4667 0.6103 0.4489 0.6312 0.4755 0.6280 0.3520 0.6381 0.3747 0.6421

### v2

@@ 3.2. Notation and Problem Setting
We divide the examinees in $\mathcal{D}_{m}$ into three attributes, $\mathcal{D}_{m}^{\text{A}},\ \mathcal{D}_{m}^{\text{B}},\ \mathcal{D}_{m}^{\text{C}}$, according to their proportion of correct responses. These groups reflect different distributions in the meta set. Let $\mathcal{A}=\{\text{A},\text{B},\text{C}\}$ denote the set of attribute categories.
  • A: the proportion of correct responses lies in $[0,0.4]$, where incorrect responses dominate.
  • B: the proportion of correct responses lies in $(0.4,0.6]$, where correct and incorrect responses are approximately balanced.
  • C: the proportion of correct responses lies in $(0.6,1.0]$, where correct responses dominate.
We also define the label set as $\mathcal{Y}=\{0,1\}$, where 0 denotes an incorrect response and 11 denotes a correct response. A group is defined as a unique combination of an attribute $a\in\mathcal{A}$ and a label $y\in\mathcal{Y}$. Based on this definition, interactions can be categorized into three types:
  • Bias-aligned samples: interactions that follow the dominant attribute pattern and are typically easier to learn, such as $\mathcal{D}_{m}^{\text{A}}$ with label 0 or $\mathcal{D}_{m}^{\text{C}}$ with label 11.
  • Unbiased samples: interactions from $\mathcal{D}_{m}^{\text{B}}$, where responses are relatively balanced and do not exhibit strong bias. These are considered neutral.
  • Bias-conflicting samples: interactions that diverge from the dominant attribute pattern and are harder to learn, for example, $\mathcal{D}_{m}^{\text{A}}$ with label 11 or $\mathcal{D}_{m}^{\text{C}}$ with label 0.
In CAT, an examinee’s true proficiency $\theta_{i}^{*}$ is assumed to remain constant, and traditional methods aim to make the final estimate $\theta_{i}^{T}$ closely approximate it. In contrast, our work focuses on mitigating selection bias in the data-driven Question Selection Module, producing more accurate and unbiased proficiency estimates while maintaining robustness under distributional shifts.
@@ 4. Method
This section introduces our approach to mitigating bias in data-driven CAT. The method targets selection bias in the meta set through a selective Mixup strategy guided by examinee attributes. By generating augmented samples from balanced response profiles, it enriches underrepresented cases and enables the Question Selection Module to learn smoother and less biased decision boundaries. The approach is simple yet effective and can be seamlessly integrated into data-driven question selection in CAT. In addition, it requires only minor modifications to existing training pipelines, making it computationally efficient and easy to implement. These properties ensure that the proposed method can be readily adopted in practical educational assessment systems.
@@ 4.1. Overview
Our method mitigates selection bias in CAT through a two-stage framework: Cross-Attribute Examinee Retrieval and Selective Mixup Regularization, as illustrated in Figure 3. In the first stage, for each examinee in $\mathcal{D}_{m}^{\text{A}}$ or $\mathcal{D}_{m}^{\text{C}}$, the most similar counterpart is retrieved from $\mathcal{D}_{m}^{\text{B}}$ based on response patterns. In the second stage, the focus is on bias-conflicting samples, such as correct responses from $\mathcal{D}_{m}^{\text{A}}$ or incorrect responses from $\mathcal{D}_{m}^{\text{C}}$. To enrich these underrepresented cases, each examinee from $\mathcal{D}_{m}^{\text{B}}$ is paired with a counterpart from $\mathcal{D}_{m}^{\text{A}}$ or $\mathcal{D}_{m}^{\text{C}}$, and Mixup is applied only when their interaction labels are identical. The final training objective combines empirical loss on original data with regularization from synthetic samples, thereby enhancing the robustness and generalization of the Question Selection Module under biased meta distributions. In the following sections, we focus primarily on the outer optimization, where these debiasing strategies are applied.
@@ 4.2. Cross-Attribute Examinee Retrieval
In real-world CAT scenarios, training data often suffers from selection bias: examinees in $\mathcal{D}_{m}^{\text{A}}$ are more likely to respond incorrectly, while those in $\mathcal{D}_{m}^{\text{C}}$ tend to respond correctly. This leads to a dominance of bias-aligned interactions, which are easy to fit but prone to shortcut learning and reduced robustness. In contrast, bias-conflicting cases (e.g., correct responses from $\mathcal{D}_{m}^{\text{A}}$ or incorrect responses from $\mathcal{D}_{m}^{\text{C}}$) are rare yet crucial for achieving fairness and generalization. To mitigate this imbalance, we leverage examinees from $\mathcal{D}_{m}^{\text{B}}$, whose balanced response patterns serve as reliable references for augmenting underrepresented cases, thereby motivating our selective Mixup design. Concretely, for each examinee $e_{a}$ with attribute $a\in\{\text{A},\text{C}\}$, we retrieve the most similar examinee $e_{b}$ with attribute $b\in\text{B}$ based on their cognitive proficiency vectors $\boldsymbol{\theta}^{*}$, estimated by the CDM. The similarity between examinees is measured using the standard $L_{2}$ norm distance:
(7) $\text{similarity}_{i,j}=\left\|\boldsymbol{\theta}_{i}^{*}-\boldsymbol{\theta}_{j}^{*}\right\|_{2}$
where $\boldsymbol{\theta}_{i}^{*}$ and $\boldsymbol{\theta}_{j}^{*}$ denote the latent knowledge proficiency vectors of examinees $i$ and $j$, respectively.
This retrieval process serves two purposes: first, it anchors bias-conflicting interactions with semantically similar but statistically more balanced samples; and second, it provides the foundation for subsequent augmentation. By aligning examinees in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$ with appropriate references from $\mathcal{D}_{m}^{\text{B}}$, the method effectively bridges biased extremes and facilitates smoother decision boundary learning in the next stage.
@@ 4.3. Selective Mixup-based Regularization
To further mitigate selection bias, we propose a selective Mixup augmentation strategy (zhang2018mixup). Unlike conventional training that relies solely on naturally occurring samples, Mixup generates synthetic interactions by interpolating instances, thereby substantially enriching the diversity of bias-conflicting data. This is particularly important in CAT, where the scarcity of such cases often causes the Question Selection Module to overfit bias-aligned patterns and neglect minority interactions. In this context, Mixup offers several advantages: it smooths decision boundaries, reduces overfitting, and improves generalization when test distributions differ from training distributions. Moreover, by diluting the influence of spurious correlations in the original data, Mixup encourages the model to learn more balanced and robust representations. Through this augmentation, the Question Selection Module is guided toward fairer and more reliable predictions, enhancing robustness and generalization under distribution shifts.
@@
Given $q_{i}\in\{\mathcal{D}_{m}^{\text{A}},\mathcal{D}_{m}^{\text{C}}\}$, its most similar counterpart in terms of proficiency is first retrieved from $\mathcal{D}_{m}^{\text{B}}$ in the previous stage. From this matched examinee in $\mathcal{D}_{m}^{\text{B}}$, an interaction $q_{j}$ with the same true label $y$ is sampled and paired with $q_{i}$, upon which synthetic bias-conflicting samples are generated through convex combination. With such paired questions $(q_{i},q_{j})$, we interpolate their parameters to generate diverse synthetic questions and enrich the training distribution, forming the synthetic set $\mathcal{D}_{\text{syn}}$. The resulting representation Mixup is defined as weighted interpolation of key parameters, where $\lambda\sim\text{Beta}(\alpha,\alpha)$ with $\alpha\in[0,1]$ controlling intensity, as shown below:
(8a) $\boldsymbol{Q}_{i,j}$ $=\lambda\boldsymbol{Q}_{i}+(1-\lambda)\boldsymbol{Q}_{j},$
(8b) $\boldsymbol{h}_{i,j}^{\text{diff}}$ $=\lambda\boldsymbol{h}^{\text{diff}}_{i}+(1-\lambda)\boldsymbol{h}_{j}^{\text{diff}},$
(8c) $h_{i,j}^{\text{disc}}$ $=\lambda h_{i}^{\text{disc}}+(1-\lambda)h_{j}^{\text{disc}},$
@@
The synthesized features are fed into the CDM, which computes the corresponding response probability as $M(q_{i,j};\boldsymbol{\theta}_{i}^{*},\zeta_{q_{i,j}},\zeta_{n})$. Since Mixup is applied only to samples with the same label, the synthetic sample keeps that label. The loss for these augmented samples is defined as:
(9) $\mathcal{L}_{\text{syn}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{q_{i,j}\in\mathcal{D}_{syn}}\ell\!\left(y_{i,q_{i,j}},M\!\left(q_{i,j};\boldsymbol{\theta}_{i}^{*},\zeta_{q_{i,j}},\zeta_{n}\right)\right),$
@@
The final outer-level optimization objective, extending Eq. (6), is expressed as a combination of empirical risk minimization and synthetic sample regularization:
(10) $\mathcal{L}_{\text{emp}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{q_{j}\in\mathcal{D}_{m}}\ell\!\left(y_{i,q_{j}},M\!\left(q_{j};\boldsymbol{\theta}_{i}^{*},\zeta_{q_{j}},\zeta_{n}\right)\right),$
(11) $\min_{\zeta_{s}}\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{emp}}+\omega\mathcal{L}_{\text{syn}},$
where $\omega$ is a regularization coefficient that balances authentic and synthetic samples.
The dual-objective design encourages the model to preserve strong predictive performance on authentic data while enhancing robustness to bias-conflicting cases. In the outer-level optimization, the Question Selection Module parameters $\zeta_{s}$ are updated via gradient descent on the combined loss. A complete overview of the training process, incorporating Mixup-based augmentation and bias-aware optimization, is provided in Algorithm 1, which summarizes the key steps and illustrates how these components are integrated into the overall framework.
Algorithm 1 Training Process of Debiased CAT
1: Initialize examinee proficiency parameters $\boldsymbol{\theta}$ in the Cognitive Diagnosis Module with pre-trained values $(\zeta_{q},\zeta_{n})$, and initialize $\zeta_{s}$ for the Question Selection Module. Specify learning rates $\alpha$ and $\beta$, the number of gradient descent steps $K$ for inner-level optimization, and the number of test steps $T$.
@@
9:   Apply selective Mixup regularization using Eqs. (8) and (9), generating synthetic bias-conflicting samples that enrich the training distribution.
@@
13:return Estimated proficiencies $\{\boldsymbol{\theta}_{i}^{*}\}_{i=1}^{N}$ and debiased Question Selection Module $\pi$.
@@
By interpolating between samples from attribute A or C and their corresponding references from attribute B, our approach introduces balanced, label-consistent synthetic data into the training process. This prevents the Question Selection Module from overfitting to dominant response patterns and encourages it to capture more nuanced relationships between question difficulty and proficiency. As a result, the model becomes less reliant on biased historical distributions, thereby improving robustness and mitigating bias across diverse examinees.
@@ 5. Experiments
In this section, we first introduce the dataset and experimental setup. Then, we conduct extensive experiments on real-world benchmark datasets. The code used in our experiments is publicly available at: https://github.com/xiaoluobouu/MDCAT.
@@ 5.1. Experimental Settings
In this section, we present the datasets, selected baselines, and the application of CAT. All models are implemented in PyTorch and trained on an NVIDIA RTX 4090 GPU to ensure efficient and consistent experimentation.
Dataset Description. We conduct experiments on two widely used educational datasets. The ASSIST0910 dataset [^2] https://sites.google.com/site/assistmentsdata/home/assistment2009-2010-data contains examinee practice logs from the ASSISTments online tutoring system, focusing on mathematics and knowledge concept problems. The NIPS-EDU dataset [^3] https://eedi.com/projects/neurips-education-challenge, released in the NeurIPS 2020 Education Challenge, consists of examinee responses collected from the Eedi learning platform, where we use the interaction data from Task 3 and Task 4. For ASSIST0910, we remove records with missing values and exclude examinees with fewer than 40 interactions. The dataset statistics, including the number of examinees, questions, concepts, and interactions, are summarized in Table 1.
@@
Partition. We randomly split examinees into training (60%), validation (20%), and testing (20%) sets. For evaluation, we perform 5-fold validation on each dataset, with each round using a different random seed. Within each round, the questions answered by each examinee are further divided into a support set ($\mathcal{D}_{u}$, 80%) and a meta set ($\mathcal{D}_{m}$, 20%), and to mitigate overfitting, these partitions are regenerated at the beginning of each training epoch. In the independent and identically distributed (IID) setting, the test split is also sampled randomly. In contrast, the out-of-distribution (OOD) setting follows a different partition strategy, which is described in detail in Section 5.2.
Table 2. Performance of different CAT models under IRT- and NCDM-based cognitive diagnosis on the unbiased ASSIST0910 and NIPS-EDU datasets. Results are reported for Metrics@5 (top block) and Metrics@10 (bottom block). Worst refers to worst-group accuracy, and Avg. refers to overall average accuracy.
@@
Metrics Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg.
Metrics@5
ERM 0.3824 0.6118 0.3400 0.6047 0.3662 0.6007 0.3268 0.5992 0.4068 0.6383 0.5064 0.6353 0.3618 0.6339 0.4479 0.6439
IRM 0.3489 0.6045 0.3280 0.5964 0.2983 0.5883 0.2860 0.5813 0.4140 0.6376 0.4450 0.6429 0.2916 0.6268 0.4098 0.6456
GroupDRO 0.5077 0.6303 0.4057 0.6134 0.3279 0.5949 0.3066 0.5894 0.3968 0.6384 0.4512 0.6427 0.2312 0.6241 0.4151 0.6432
Reweight 0.5328 0.6395 0.3967 0.6142 0.4793 0.6152 0.4362 0.6107 0.5699 0.6363 0.4832 0.6272 0.4040 0.6374 0.4172 0.6468
Ours 0.5470 0.6393 0.4382 0.6178 0.4989 0.6116 0.4823 0.6124 0.5701 0.6361 0.5185 0.6371 0.4250 0.6386 0.4315 0.6473
Metrics@10
@@
Ours 0.4489 0.6320 0.3854 0.6095 0.4426 0.6117 0.4753 0.6155 0.4489 0.6312 0.4674 0.6268 0.3520 0.6381 0.3517 0.6436
Evaluation. We evaluate all methods on their ability to predict binary-valued responses in the meta set $\mathcal{D}_{m}$, using two complementary metrics: worst-group accuracy (Worst) and overall average accuracy (Avg.). The Worst metric reflects model robustness by focusing on the most disadvantaged groups, while the Avg. metric captures the overall prediction performance across all examinees.
Baseline Methods. For CDMs, we adopt the 1PL IRT model (embretson2013irt) and NCDM (wang2020ncd), both implemented in the EduStudio [^4] https://edustudio.ai/ framework (wu2025edustudio). For CAT frameworks, we consider data-driven approaches that formulate CAT as a bi-level optimization problem solved via meta-learning, including BOBCAT (ghosh2021bobcat) and UATS (yu2024ucats). BOBCAT introduces a learnable Question Selection Module, whereas UATS employs a differentiable hierarchical optimization framework to improve efficiency and adaptability. To better analyze bias in question selection, we freeze the question parameters of the CDM during outer-loop optimization and initialize them with pre-trained values.
To validate the effectiveness of our method, we compare it against the following debiasing baselines, all implemented within the same bi-level training framework and adapted to the CAT setting to ensure consistent evaluation across different situations.
ERM (Empirical Risk Minimization) minimizes the average prediction loss over the entire meta set without explicitly addressing distributional bias. It serves as a standard baseline but often overfits to dominant bias-aligned patterns.
IRM (arjovsky2019irm) (Invariant Risk Minimization) encourages feature representations that yield consistent predictions across environments. In our setting, we simulate multiple environments by sampling examinees with different attribute values (A, B, and C). This approach ensures that the Question Selection Module performs robustly across these diverse distributions.
GroupDRO (sagawa2019groupDRO) (Group Distributionally Robust Optimization) improves worst-case performance across predefined groups. It optimizes the Question Selection Module by minimizing the maximum group-wise loss. This strategy helps prevent collapse on underrepresented or difficult subpopulations.
Reweight (japkowicz2000reweight) mitigates imbalance by assigning larger weights to bias-conflicting samples during training. This guides the model to better capture underrepresented patterns and reduces reliance on shortcut signals from bias-aligned data.
Parameter Settings. In NCDM, the two hidden layers contain 128 and 64 neurons, which control the network’s capacity and complexity and allow it to capture nonlinear relationships between examinees and items. The $\alpha$ parameter of the Beta distribution is chosen from $\{0.2,0.4,0.6,0.8,1.0\}$, and the mixing coefficient $\omega$ in the Mixup loss is selected from the same range to ensure consistent regularization strength. To approximate each examinee’s proficiency, inner optimization updates the latent proficiency vector $K$ times, with $K=5$ in our experiments to balance accuracy and computational cost.